OpenStudy (anonymous):

]If f(x) = x2 + 2, find f(4). 10 18 2 6

OpenStudy (anonymous):

f(x) = x2 + 2 just put 4 every where you see x

OpenStudy (anonymous):

f(4)=(4)(2)+2 f(4)=8+2 f(4)=10